OpenStudy (anonymous):

Determine the line of intersection of each system of equations. help would be greatly appreciated :D a.) 2x + y + z = 7 4x +3y - 3z = 13 4x _2y +2z = 14

hero (hero):

1.) 2x + y + z = 7 2.) 4x +3y - 3z = 13 3.) 4x - 2y +2z = 14 Subtract Equation 3 from Equation 2 to get the line. Then subtract Equation 1 from Equation 2 to verify. You want to eliminate x variable in both subtractions.

OpenStudy (texaschic101):

2x + y + z = 7 -->(-2)2x + y + z = 7 4x + 3y - 3z = 13 ----------------- -4x - 2y - 2z = - 14 (result of multiplying by -2) 4x + 3y - 3z = 13 -----------------add y - 5z = -1 4x + 3y - 3z = 13 -->(-1)4x + 3y - 3z = 13 4x - 2y + 2z = 14 ---------------- -4x - 3y + 3z = - 13 (result of multiplying by -1) 4x - 2y + 2z = 14 -----------------add - 5y + 5z = 1 y - 5z = -1 -5y + 5z = 1 --------------add -4y = 0 y = 0/-4 = 0 y - 5z = - 1 0 - 5z = - 1 -5z = - 1 z = -1/-5 = 1/5 2x + y + z = 7 2x + 0 + 1/5 = 7 2x = 7 - 1/5 2x = 35/5- 1/5 2x = 34/5 x = (34/5)/2 x = 34/5 * 1/2 x = 34/10 = 17/5 check... 2x + y + z = 7 2(17/5) + 0 + 1/5 = 7 34/5 + 0 + 1/5 = 7 35/5 = 7 7 = 7 (correct) check... 4x + 3y - 3z = 13 4(17/5) + 3(0) - 3(1/5) = 13 68/5 + 0 - 3/5 = 13 65/5 = 13 13 = 13 (correct) check... 4x - 2y + 2z = 14 4(17/5) - 2(0) + 2(1/5) = 14 68/5 - 0 + 2/5 = 14 70/5 = 14 14 = 14 (correct) ANSWER : x = 17/5, y = 0, z = 1/5 or (17/5,0,1/5)

hero (hero):

@texaschic101 : It says to determine the `line of intersection`, not the points of intersection.

OpenStudy (texaschic101):

oops

hero (hero):

I think this question should have read: If possible, determine the line of intersection of each system of equations.